The number of degrees in an acute angle of a right angled triangle is equal to the number of grades in other ;express both the angles in degrees

Grade:11

1 Answers

Vikas TU
6 years ago
 we Know that, 
180 degree = 200 review 
so 1 review = 9/10 degree 
let 
x degree + x review =90 degree 
x degree + 9x/10 degree = 90 degree 
19x/10 = 90 degree 
x = 900/19 
so edges are 900/19 degree and 810/19 degree